OpenStudy (anonymous):

sin x/1-cos x + sin x/1+cos x

OpenStudy (john_es):

Do the mcm of the fractions, and sum. \[\frac{\sin x(1+\cos x)+\sin x(1-\cos x)}{1-\cos^2x}\] Follow from this point.

OpenStudy (anonymous):

I have that but from there im lost